# Pontryagin Optimal Control via Neural Networks

### 07 Jan 2024 v2

@@ Abstract
Solving real-world optimal control problems are challenging tasks, as the complex, high-dimensional system dynamics are usually unrevealed to the decision maker. It is thus hard to find the optimal control actions numerically. To deal with such modeling and computation challenges, in this paper, we integrate Neural Networks with the Pontryagin’s Maximum Principle (PMP), and propose a sample efficient framework NN-PMP-Gradient. The resulting controller can be implemented for systems with unknown and complex dynamics. By taking an iterative approach, the proposed framework not only utilizes the accurate surrogate models parameterized by neural networks, it also efficiently recovers the optimality conditions along with the optimal action sequences via PMP conditions. Numerical simulations on Linear Quadratic Regulator, energy arbitrage of grid-connected lossy battery, control of single pendulum, and two MuJoCo locomotion tasks demonstrate our proposed NN-PMP-Gradient is a general and versatile computation tool for finding optimal solutions. And compared with the widely applied model-free and model-based reinforcement learning (RL) algorithms, our NN-PMP-Gradient achieves higher sample-efficiency and performance in terms of control objectives.
@@ I Introduction
Solving real-world optimal control problems is possessing huge engineering and societal values, ranging from running power grids, achieving robotics manipulation, and operating transportation networks. However, finding such real-world optimal controllers is often a challenging task, since real-world systems dynamics are usually unrevealed and complex due to their uniqueness, non-linearity and high-dimension [1, 2]. Previous literature provides several extensively utilized approaches on building optimal controllers, including dynamic programming (DP) and model predictive control (MPC) [3]. However, all of these approaches require full knowledge of underlying state transitions. Such restrictions address difficulties on directly applying these approaches in solving complicated real-world optimal control problems with unknown system dynamics [4, 5].
Recent developments of Reinforcement Learning (RL) provides another direction to tackle this problem. Unlike previous model-based approaches, deep RL algorithms relax the requirements over prior knowledge about system dynamics [6]. Instead, agents can be trained to either learn the state-action value function (Q function) or learning the policies directly. Hence, these RL algorithms are termed as model-free approaches [7]. Typical deep RL methods achieve impressive performances on certain control tasks [8, 9]. However, model-free deep RL methods often suffer from sample inefficiency. They usually require huge amount of trials to achieve effective solution, leading to considerable training time, which may also be impractical for real-world applications. Moreover, in general current model-free RL lacks interpretability  [10], which raises reliability and safety concerns, especially for problems with hard constraints [11, 12].
To address these challenges, researchers revisit model-based approaches and develop model-based reinforcement learning (MBRL). Such learning-based method usually first acquires a model of system dynamics and then find control policies based on learned models [13]. Recent introduction of deep neural networks (NNs) promises a versatile and efficient modeler for complex, high-dimensional system dynamics [14]. Such surrogate learned models are deployed in model-based RL (MBRL) frameworks [15, 16]. However, typical MBRL algorithms such as MPC with random shooting (RS-MPC) [17] or learning-based policy networks do not provide an explicit mechanism for finding optimal decisions [18]. The lack of tractability severely influence the control performance and reliability of these MBRL algorithms, hence restricting their potential on real-world optimal control problems with unknown dynamics.
@@
In this paper, we design a tractable, sample-efficient MBRL with accurate learnt dynamics model. We start from the idea of learning complex, unknown system dynamics with neural networks (NNs) in MBRL. More importantly, we leverage the potential of Pontryagin’s maximum principle (PMP), which offers a theoretical foundation for optimality conditions, and has been used to numerically solve the optimal control actions for a few decades [19]. PMP states a set of necessary conditions, including a minimum condition of the control Hamiltonian, for optimal control policies and optimal state trajectory. Unlike DP or MPC, solving PMP conditions can give the entire optimal solutions over the planning horizon. Working examples on control of electric vehicles [20], supply chain [21], and financial markets [22] have been brought up. Several recent works also considered data-driven design of PMP-based controllers. [23, 24] introduce an auxiliary system by differentiating PMP conditions of given systems, to help take gradients for iteratively updating the parameters of a policy network. In [25] authors directly apply PMP conditions and the Hamiltonian-Jacobian-Bellman (HJB) equations to guide the policy network training. Yet these approaches still require computationally expensive training and solution processes.
Different from previous design of using policy network as controller, our major algorithm developed in this work, named NN-PMP-Gradient, directly establishes a differentiable algorithm PMP-Gradient over PMP conditions to numerically solve the optimal control actions with learned NN system dynamics. We also prove that with accurate NN model, PMP-Gradient controller can at least ensure convergence to local optimal solution, and under certain conditions can converge to global optimal. Numerical simulations on a variety of control tasks including Linear Quadratic Regulator, charging of grid-connected lossy battery, control of single pendulum, and two MuJoCo locomotion, convincingly demonstrate our proposed NN-PMP-Gradient can provide a feasible optimal control. The results also show that our NN-PMP-Gradient: (1) models complex, unknown system dynamics more accurately than traditional Linear models, and hence achieves better control performance. (2) has higher sample efficiency and provide safer solution within state bounds. (3) usually has better control performance than popular model-free RL method PPO, classical model-based RL method RS-MPC, under same learnt NN-dynamics.
@@ III Model-based Optimal Controller Design
In this section, we demonstrate our proposed model-based approach for optimal control named NN-PMP-Gradient, which is composed of two major parts: i). train neural networks (NN) to learn system dynamics offline from data; and ii). use a PMP-based gradient method to obtain the optimal control policy based on the learned dynamics. Proposed scheme of NN-PMP-Gradient is illustrated in Fig. 1.
@@ III-A System Identification with Neural Networks
We would like to leverage PMP’s advantages on finding optimal control sequences, while the system dynamics are unknown. To realize it, we first need an effective method to identify the unrevealed system dynamics, since applying PMP requires the explicit expression of system dynamics. In our proposed approach, we utilize the function approximation capabilities of NN, and parameterize the system dynamics either wholly or partially as a neural network:
$\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t})\quad\text{with given $\mathbf{x}_{0}$};$ (7)
Usually we train the NN-simulated system dynamics $f_{NN}$ offline with dataset $\mathbb{D}=\{\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)}|\mathbf{x}_{t+1}^{(i)}\}_{i=1,2,...,N}$. To establish $\mathbb{D}$, we randomly collected state-control pairs $\{\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)}\}$ from state and control space $\mathcal{X}$, $\mathcal{U}$ as features, and use ground-truth system dynamics $f$ to generate next state $\mathbf{x}_{t+1}^{(i)}$ as supervised labels. Once the training data is collected, we use mean squared error (MSE) to train $f_{NN}$:
$\mathcal{L}_{sysid}=\frac{1}{N}\sum_{i=1}^{N}(f_{NN}(\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)})-\mathbf{x}_{t+1}^{(i)})^{2}.$ (8)
To improve the dynamics approximation accuracy, we can also train $f_{NN}$ in an online fashion by updating $\mathbb{D}$ with new state-control pairs executed in the control process and re-training $f_{NN}$, which is also broadly used in the Reinforcement Learning community. Based on the trained system dynamics and designated optimizing objective, we formulate the Hamiltonian and establish equations for optimal control and state trajectory with PMP Conditions Eq. (4)-(6). However, as non-linearity shows up in NN-simulated system dynamics $f_{NN}$, we still face challenges in directly solving these equations, which motivates us to find numerical way to solve these equations with learned dynamics involved.
@@ III-B PMP-Gradient Controller Design
With dynamics accurately parameterized by the learned function $f_{NN}$ in Eq. (7), we are able to establish PMP conditions Eq. (4)-(6) for the system. While the ultimate goal is to find solutions for these equations, which help us get the optimal control policy $\pi=\mathbf{u}_{0:T-1}^{*}$ of the sequential decision-making problems. Due to the non-linearity and complicated function space of $f_{NN}$, it is challenging to directly obtain solutions in an analytical way. To overcome the difficulty, we design a gradient-based method named PMP-Gradient, to compute the numerical value of $\pi$.
In this approach, we start from an initialized random control policy $\mathbf{u}^{(1)}_{0:T-1}$, and use it to retrieve corresponding state sequence $\mathbf{x}^{(1)}_{0:T}$ based on NN-simulated system dynamics Eq. (7). Then based on it, we recursively compute costate $\mathbf{\lambda}^{(1)}_{0:T}$ with PMP Conditions Eq. (4)-(5):
@@
Algorithm 1 PMP-Gradient Controller
1:Initial state $\mathbf{x}_{0}$; Maximum iterations $M$; Learning rate $\eta$; NN-simulated system dynamics $\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t})$;
@@
5:while $k\leq M$ do
6:     Compute $\mathbf{x}_{0:T}^{(k)}$ with current control policy $\mathbf{u}_{0:T-1}^{(k)}$, based on $\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t})$ and $\mathbf{x}_{0}$.
7:     Compute terminal costate $\mathbf{\lambda}_{T}^{(k)}$ by $\mathbf{\lambda}_{T}^{(k)}=\frac{\partial\Phi(\mathbf{x}_{T}^{(k)})}{\partial\mathbf{x}_{T}^{(k)}}$.
8:     Compute costate $\mathbf{\lambda}_{t}^{(k)}$ for $t=0,1,..,T-1$, based on $\mathbf{\lambda}_{t}^{(k)}=\frac{\partial H(\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)},\mathbf{\lambda}_{t+1}^{(k)})}{\partial\mathbf{x}_{t}^{(k)}}$ and $\mathbf{\lambda}_{T}^{(k)}$.
9:     Update Control Policy: $\mathbf{u}_{t}^{(k+1)}\leftarrow\mathbf{u}_{t}^{(k+1)}-\eta\frac{\partial H(\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)},\mathbf{\lambda}_{t+1}^{(k)})}{\partial\mathbf{u}_{t}^{(k)}}$.
10:     $k\leftarrow k+1$
11:end while
12:$\mathbf{u}_{0:T-1}^{*}\leftarrow\mathbf{u}_{0:T-1}^{(k)}$
@@
Essentially, we utilize a fully differentiable procedure by using the learned dynamics $f_{NN}$ with the corresponding Hamiltonian and the terminal cost. When $f_{NN}$ approximates ground-truth $f$ accurately, we have the following Lemma regarding step $5-6$ of PMP-Gradient Algorithm 1 and Theorem for cost $J$ in optimal control problem Eq. (2):
@@
Based on Lemma 1 and Theorem 2, we can infer that our PMP-Gradient Controller has global convergence guarantees with convergence rate $\mathit{O}(1/\eta)$ when $J$ is a convex function w.r.t. control variable $\mathbf{u}_{t}$ for all $t=0,1,..,T-1$ (and has at least local convergence guarantees for general cases). The proof of Lemma 1 and Theorem 2 can be found in Appendix. -A in our paper’s online version [27]. To accelerate convergence process, we can also apply other gradient-based optimizer like conjugate gradient (CG) optimizer, Newton’s method to replace gradient descent in step 7 in Algorithm 1.
Different from directly performing gradient descent over objective $J$, our PMP-Gradient controller has no need to consider terms with future states $\mathbf{x}_{t+1:T}$ when taking derivatives of $H(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{\lambda}_{t+1})$ w.r.t. $\mathbf{u}_{t}$. This helps us eliminate computations of Hessian matrices $\frac{\partial\mathbf{x}_{t+1}}{\partial\mathbf{x}_{t}}$, hence reducing the size of computation graph for calculating gradients and greatly improving the computation time. For problems with hard state and action bounds, we modify objective and step 7 in Algorithm 1. Detailed settings can be found in Appendix B in our paper’s online version [27].

### 15 Jan 2024 v3

@@ Abstract
Solving real-world optimal control problems are challenging tasks, as the complex, high-dimensional system dynamics are usually unrevealed to the decision maker. It is thus hard to find the optimal control actions numerically. To deal with such modeling and computation challenges, in this paper, we integrate Neural Networks with the Pontryagin’s Maximum Principle (PMP), and propose a sample efficient framework Neural-PMP. The resulting controller can be implemented for systems with unknown and complex dynamics. By taking an iterative approach, the proposed framework not only utilizes the accurate surrogate models parameterized by neural networks, it also efficiently recovers the optimality conditions along with the optimal action sequences via PMP conditions. Numerical simulations on Linear Quadratic Regulator, energy arbitrage of grid-connected lossy battery, control of single pendulum, and two MuJoCo locomotion tasks demonstrate our proposed Neural-PMP is a general and versatile computation tool for finding optimal solutions. And compared with the widely applied model-free and model-based reinforcement learning (RL) algorithms, our Neural-PMP achieves higher sample-efficiency and performance in terms of control objectives.
@@ I Introduction
Solving real-world optimal control problems is possessing huge engineering and societal values, ranging from running power grids, achieving robotics manipulation, and operating transportation networks. However, finding such real-world optimal controllers is often a challenging task, since real-world systems dynamics are usually unrevealed and complex due to their uniqueness, non-linearity and high-dimension [1, 2]. Previous literature provides several extensively utilized approaches on building optimal controllers, including dynamic programming (DP) and model predictive control (MPC) [3]. However, most of these approaches require full knowledge of underlying state transitions, or use reduced-order models to facilitate numerical solution process for higher-dimensional problems. Such restrictions address difficulties on directly applying these approaches in solving complicated real-world optimal control problems with unknown system dynamics [4, 5].
Recent developments of Reinforcement Learning (RL) provides another direction to tackle this problem. Model-free deep RL algorithms relax the requirements over prior knowledge on system dynamics [6]. Instead, agents can be trained to either learn the state-action value function (Q function) or learning the policies directly. Typical deep RL methods achieve impressive performances on certain control tasks [7, 8]. However, model-free deep RL methods often suffer from sample efficiency, requiring huge amount of trials and training time to fit effective policy networks. Moreover, typical model-free RL lacks interpretability  [9], which raises reliability and safety concerns, especially for problems with hard constraints [10, 11]. To address these challenges, researchers revisit model-based approaches and develop model-based RL (MBRL). Such learning-based method usually first acquires a model of system dynamics and then find control policies based on learned models [12]. Recent introduction of deep neural networks (NN) in data-driven control promises a versatile and efficient tool for controlling under unknown dynamics [13]. Such surrogate learned models are deployed in model-based RL (MBRL) frameworks [14, 15]. Compared with model-free RL with only learning-based policy networks, NN dynamics in MBRL exhibit higher generalizability and effectiveness in unexplored state-action regions. The generalizability of NN dynamics helps significantly reducing the training samples for MBRL methods. Yet typical MBRL algorithms such as MPC with random shooting (RS-MPC) [16] or policy networks do not provide explicit mechanism for finding optimal decisions [17]. The lack of tractability severely influence the control performance and reliability of these MBRL algorithms, hence restricting their potential on real-world problems with unknown dynamics.
@@
In this paper, we design a novel and sample-efficient controller built upon accurate dynamics purely learned from data as well as theoretical foundations of Pontryagin’s maximum principle (PMP). We leverage the potential of PMP into data-driven controller design with a fully differentiable architecture. PMP has been used to numerically solve the optimal control problems for a few decades [18]. It states a set of necessary conditions, including a minimum condition of the control Hamiltonian, for optimal control policies and optimal state trajectory. Unlike DP or MPC, solving PMP conditions can give the entire optimal solutions over the planning horizon. Working examples on control of electric vehicles [19], supply chain [20], and financial markets [21] have been brought up. However, PMP’s effectiveness relies on knowing the exact control Hamiltonian, while it is less investigated for systems with unknown or unrevealed dynamics. To overcome such challenge, we seamlessly integrate learning dynamics purely from collected state-action samples with NN into PMP-based optimal control. Once NN dynamics are well learned, we explicitly write out the Hamiltonian parameterized by the NN, making it possible to solve the derived optimality conditions using the surrogate NN model.
Several recent works also considered data-driven design based on PMP principles. [22, 23] introduce auxiliary models by differentiating PMP conditions, while the major objective for these methods are to help take gradients for policy network updates. In [24] authors directly apply PMP conditions and the Hamiltonian-Jacobian-Bellman (HJB) equations to guide the policy network training. Yet these approaches still involve sampling-expensive policy network fitting processes, and does not provide guarantee on the tractability of control actions. Different from previous design of using policy network as controller, our major algorithm, Neural-PMP, directly establishes a differentiable procedure PMP-Gradient over PMP conditions to numerically solve the optimal control actions with learned NN system dynamics. Such iterative approach has been discussed in the literature, such as the continuous-time steepest gradient approach introduced in [25] for solving equations with PMP conditions involved. Neural-PMP differentiates from these methods in greatly expanding the applicability of PMP conditions under the scenarios with only data and NN model available. We also prove with accurate NN model, PMP-Gradient controller can at least ensure convergence to local optimal solution, and under certain conditions can converge to global optimal. Numerical simulations on a variety of control tasks including Linear Quadratic Regulator, charging of grid-connected lossy battery, control of single pendulum, and two MuJoCo locomotion tasks convincingly demonstrate our proposed Neural-PMP can: (1) outperform standard Linearized MPC models in control performance thanks to more accurate NN modeling; (2) achieve superior control performance with fewer training samples compared to deep RL algorithms, showcasing higher sample efficiency; (3) promote safer solutions with regard to state and action bounds versus MBRL using random search or model-free RL.
@@ III Model-based Optimal Controller Design
In this section, we demonstrate our proposed model-based approach for optimal control named Neural-PMP, which is composed of two major parts: i). train neural networks (NN) to learn system dynamics offline from data; and ii). use a PMP-based gradient method to obtain the optimal control policy based on the learned dynamics. Proposed scheme of Neural-PMP is illustrated in Fig. 1.
@@ III-A System Identification with Neural Networks
To leverage PMP’s advantages on finding optimal control sequences of unknown dynamics, we first need an effective method to model such dynamics from collected data. By doing so, we are able to formulate the required control Hamiltonian explicitly. In our proposed approach, we utilize the function approximation capabilities of NN, and parameterize the system dynamics either wholly or partially as NN:
$\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t};\theta)\quad\text{with given $\mathbf{x}_{0}$};$ (7)
Usually we train the NN-simulated system dynamics $f_{NN}$ offline with dataset $\mathbb{D}=\{\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)}|\mathbf{x}_{t+1}^{(i)}\}_{i=1,2,...,N}$. To establish $\mathbb{D}$, we randomly collect state-control pairs $\{\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)}\}$ from state and control space $\mathcal{X}$, $\mathcal{U}$, and use ground-truth system dynamics $f$ to generate next state $\mathbf{x}_{t+1}^{(i)}$ as supervised labels. Once the training data is collected, we learn parameters $\theta$ of $f_{NN}$ by minimizing mean squared error (MSE):
$\min_{\theta}\;\frac{1}{N}\sum_{i=1}^{N}(f_{NN}(\mathbf{x}_{t}^{(i)},\mathbf{u}_{t}^{(i)};\theta)-\mathbf{x}_{t+1}^{(i)})^{2}.$ (8)
To improve the dynamics approximation accuracy, we can also train $f_{NN}$ in an online fashion by updating $\mathbb{D}$ with new state-control pairs executed in the control process and re-training $f_{NN}$, which is also broadly used in RL community. Based on the trained system dynamics and designated optimizing objective, we formulate the Hamiltonian and establish equations for optimal control and state trajectory with PMP Conditions Eq. (4)-(6). However, as non-linearity shows up in NN-simulated system dynamics $f_{NN}$, we still face challenges in directly solving these equations, which motivates us to find numerical way to solve these equations with learned dynamics involved.
@@ III-B PMP-Gradient Controller Design
With accurately learned function $f_{NN}$ in Eq. (7), we are able to establish control Hamiltonian $H(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{\lambda}_{t+1})=L(\mathbf{x}_{t},\mathbf{u}_{t})+\mathbf{\lambda}_{t+1}^{{}^{\prime}}f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t};\theta)$, as well as PMP conditions Eq. (4)-(6) for the system. While the ultimate goal is to find solutions for these equations, which help us get the optimal control policy $\pi=\mathbf{u}_{0:T-1}^{*}$ of the sequential decision-making problems. Due to the non-linearity and complicated function space of $f_{NN}$, it is challenging to directly obtain solutions in an analytical way. Algorithms such as random shooting has tractability and safety concerns, as there is no guarantee on action’s performance, while the resulting dynamics are not regulated. To overcome these difficulties, we are inspired by the iterative approaches and design gradient-based method PMP-Gradient to compute numerical solutions for $\pi$.
In this approach, we start from an initialized random control policy $\mathbf{u}^{(1)}_{0:T-1}$ with superscript denoting action updating iterations. And we use such control to retrieve corresponding state sequence $\mathbf{x}^{(1)}_{0:T}$ based on NN-simulated system dynamics Eq. (7). Then based on it, we recursively compute costate $\mathbf{\lambda}^{(1)}_{0:T}$ with PMP Conditions Eq. (4)-(5):
@@
Algorithm 1 Neural-PMP
1:Initial state $\mathbf{x}_{0}$; Maximum iterations $M$; Learning rate $\eta$; Dataset $\mathbb{D}$ for training NN-simulated system dynamics $\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t};\theta)$;
@@
5:Learn $\theta$ for NN dynamics $f_{NN}$ with $\mathbb{D}$.
6:while $k\leq M$ do
7:     Establish form of Hamiltonian: $H(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{\lambda}_{t+1})=L(\mathbf{x}_{t},\mathbf{u}_{t})+\mathbf{\lambda}_{t+1}^{{}^{\prime}}f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t};\theta)$
8:     Compute $\mathbf{x}_{0:T}^{(k)}$ with current control policy $\mathbf{u}_{0:T-1}^{(k)}$, based on $\mathbf{x}_{t+1}=f_{NN}(\mathbf{x}_{t},\mathbf{u}_{t})$ and $\mathbf{x}_{0}$.
9:     Compute terminal costate $\mathbf{\lambda}_{T}^{(k)}$ by $\mathbf{\lambda}_{T}^{(k)}=\frac{\partial\Phi(\mathbf{x}_{T}^{(k)})}{\partial\mathbf{x}_{T}^{(k)}}$.
10:     Compute costate $\mathbf{\lambda}_{t}^{(k)}$ for $t=0,1,..,T-1$, based on $\mathbf{\lambda}_{t}^{(k)}=\frac{\partial H(\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)},\mathbf{\lambda}_{t+1}^{(k)})}{\partial\mathbf{x}_{t}^{(k)}}$ and $\mathbf{\lambda}_{T}^{(k)}$.
11:     Update Control Policy: $\mathbf{u}_{t}^{(k+1)}\leftarrow\mathbf{u}_{t}^{(k)}-\eta\frac{\partial H(\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)},\mathbf{\lambda}_{t+1}^{(k)})}{\partial\mathbf{u}_{t}^{(k)}}$.
12:     (If apply online training)  Apply $\mathbf{u}_{0:T-1}^{(k)}$ to real system $f$, collect state-action pairs $\{\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)}|f(\mathbf{x}_{t}^{(k)},\mathbf{u}_{t}^{(k)})\}_{t=0:T-1}$, add them to $\mathbb{D}$ and re-train $f_{NN}$.
13:     $k\leftarrow k+1$
14:end while
15:$\mathbf{u}_{0:T-1}^{*}\leftarrow\mathbf{u}_{0:T-1}^{(k)}$
@@
Essentially, we utilize a fully differentiable procedure by using the learned dynamics $f_{NN}$ with the corresponding Hamiltonian and the terminal cost. Unlike typical policy network updates in model-free RL, the inference of control actions in our Neural-PMP is training-free, which further reduce computation need during training. Meanwhile, when $f_{NN}$ approximates ground-truth $f$ accurately, we have the following Lemma regarding step $5-6$ of Neural-PMP Algorithm 1 and Theorem for cost $J$ in w.r.t Eq. (2):
@@
Based on Lemma 1 and Theorem 2, we can infer that our PMP-Gradient Controller has global convergence guarantees with convergence rate $\mathit{O}(1/\eta)$ when $J$ is a convex function w.r.t. control variable $\mathbf{u}_{t}$ for all $t=0,1,..,T-1$ (and has at least local convergence guarantees for general cases). The proof of Lemma 1 and Theorem 2 can be found in Appendix. -A in our paper’s online version [27]. These observations demonstrate the generalizability of our Neural-PMP based on accurately trained NN dynamics across the entire state-action space. Indeed, our approach is more sample efficient as fitting dynamics is much more efficient than training policy networks in model-free RL [16]. To accelerate policy convergence, other gradient-based optimizers like conjugate gradient (CG) or Newton’s method can be also applied.
Different from directly performing gradient descent over objective $J$, PMP-Gradient controller does not need to consider terms with future states $\mathbf{x}_{t+1:T}$ when taking derivatives of $H(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{\lambda}_{t+1})$ w.r.t. $\mathbf{u}_{t}$. This helps us eliminate computations of Hessian matrices $\frac{\partial\mathbf{x}_{t+1}}{\partial\mathbf{x}_{t}}$, hence reducing the size of computation graph for calculating gradients and greatly improving the computation time. For problems with hard state and action bounds, we modify objective and step 7 in Algorithm 1. Detailed settings can be found in Appendix B in our paper’s online version [27]. With this design, we can mitigate the risk of state and action exceeding bounds, as this controller avoids any uncontrollable random elements.